OpenStudy (timaashorty):

Anybody generous to give me a few minutes of their time ?(: I need help in factoring quadratics, and I want to make sure If I'm getting it or not. -help will be greatly appreciated <3

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

It helps if you understand how to factor, a little. But if you're lazy and terrible at factoring, you can use it to get the factors. :D

OpenStudy (anonymous):

(I am lazy and terrible at factoring. So I do this.)

OpenStudy (timaashorty):

The factoring formula ? x^2+bx+c = 0 ? Or y= ax^2+bx+c ?

OpenStudy (anonymous):

No. \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]If you don't know what that is, don't pay attention to me. You'll learn it very soon.

OpenStudy (anonymous):

what's your question?

OpenStudy (timaashorty):

I'm not ready for that yet. I want to learn it this way: ax2 + bx + c = 0 first.

OpenStudy (anonymous):

Ok, you want to find two numbers that multiply to get b and add to get c.

OpenStudy (timaashorty):

My question is that I want to practice factoring, I'm beginning to get it, but I still need practice.

OpenStudy (anonymous):

so are you trying to ax^2+bx+c=0 into maybe something like (x+2)(x+6) ?

OpenStudy (anonymous):

oh factor this then (x^2)-4

OpenStudy (timaashorty):

YES PENGUIN (:

OpenStudy (timaashorty):

4x^2+11x+6 How would I factor that ?

OpenStudy (timaashorty):

Basically I'm looking for the steps, like do it together. lol

OpenStudy (anonymous):